# Incentive-Aligned Multi-Source LLM Summaries

### v1

@@ Appendix G Practical Notes and Scaling for Finite-K Guarantees
    Computational Cost: Computing the score $\widehat{w}_{i}$ for one source requires averaging over $K$ claims and $|\mathcal{C}|-1$ peers, resulting in a cost of $O(K(|\mathcal{C}|-1))$. Scoring all sources takes $O(|\mathcal{C}|K(|\mathcal{C}|-1))$. Generating the random permutation for the off-task baseline costs $O(K)$.
  4. 4.
@@
Appendix H Claim-wise Heterogeneity in Coverage, Signals, and Reporting
We extend the analysis to claim-wise heterogeneity for the scored source $i$: coverage $\alpha_{ik}$, signal quality $\eta_{i}^{\mathrm{sig}}(k)$, and reporting parameters $(q_{1k},q_{0k})$, while peers $j\neq i$ remain claim-dependent as in the main text. Effort is global and binary: if $e_{i}{=}0$ then $\eta_{i}^{\mathrm{sig}}(k){=}0$ for all $k$; if $e_{i}{=}1$ then $\eta_{i}^{\mathrm{sig}}(k)\geq 0$. Utilities are $u_{i}(F_{i})=v_{i}\,\Pr(\widehat{w}_{i}\geq t_{\mathrm{src},i})-c_{i}e_{i}$ with $v_{i}>c_{i}$.
Assumptions.
A1–A2 (independent claim blocks with common prior $\pi_{i}$, post-selection conditional independence) hold as stated. Coverage is non-anticipatory and independent across sources: $Q_{ik}=\mathbf{1}\{r_{ik}\neq\bot\}$ with $\alpha_{ik}:=\Pr(Q_{ik}=1\mid T_{i})$, and $Q_{ik}\perp Q_{jk}\mid T_{i}$. For peer margin we strengthen A3 to:
$2\pi_{i}(1-\pi_{i})\,\Gamma_{i}(k)\ \geq\ \gamma_{\min}\ >0\quad\text{for all }k,\qquad\text{where }\Gamma_{i}(k):=\mathbb{E}_{j\neq i}[\alpha_{jk}\eta_{jk}\mid T_{i}].$ (A3)
(Thus the average margin $\bar{\gamma}=\tfrac{1}{K}\sum_{k}2\pi_{i}(1-\pi_{i})\Gamma_{i}(k)\geq\gamma_{\min}>0$.) This is a stronger but still reasonable assumption when a prefilter for the RAG system yields an on-average reliable peer pool for each claim.
Signals and reporting.
Under effort, $z_{ik}\in\{0,1\}$ with $s_{1}(k)=\Pr(z_{ik}{=}1\mid\theta_{k}{=}1)$, $s_{0}(k)=\Pr(z_{ik}{=}1\mid\theta_{k}{=}0)$, and $\eta_{i}^{\mathrm{sig}}(k):=s_{1}(k)-s_{0}(k)\geq 0$. Reporting may vary by claim:
$q_{1k}=\Pr(r_{ik}{=}1\mid z_{ik}{=}1,Q_{ik}{=}1),\qquad q_{0k}=\Pr(r_{ik}{=}1\mid z_{ik}{=}0,Q_{ik}{=}1),\qquad$
$\eta_{i}(k)=(q_{1k}-q_{0k})\,\eta_{i}^{\mathrm{sig}}(k)\leq\eta_{i}^{\mathrm{sig}}(k),$
with equality when $(q_{1k},q_{0k})=(1,0)$ (claim-wise truthful reporting).
What changes vs. the main text.
All statements and proofs go through with minor changes (highlighted eblow) after replacing the claim-invariant factors $\alpha_{i}\eta_{i}$ by their claim-wise counterparts $\alpha_{ik}\eta_{i}(k)$ inside the per-claim summand and averaging over $k$. The off-task pairing and the bounded-differences constant remain the same.
Proposition 2 (Expected score with heterogeneity).
Under A1–A2 and the coverage conditions,
$\mathbb{E}[\widehat{w}_{i}\mid q,T_{i}]\ =\ \frac{1}{K}\sum_{k=1}^{K}2\pi_{i}(1-\pi_{i})\,\alpha_{ik}\,\eta_{i}(k)\,\Gamma_{i}(k).$
Proof. Identical to Proposition 3.1, substituting $\alpha_{ik}$ for $\alpha_{i}$ and $\eta_{i}(k)$ for $\eta_{i}$ inside each claim’s on-task term and in the off-task baseline before averaging over $k$.
Concentration.
Changing one claim block affects at most three per-claim terms; hence $|\Delta\widehat{w}_{i}|\leq 3/K$ as in Lemma E, and McDiarmid’s inequality (Theorem E) gives the same sub-Gaussian tail.
Asymptotic informed-truthfulness (unchanged in spirit).
Define
$\mu_{i}^{\mathrm{truth}}:=\frac{1}{K}\sum_{k=1}^{K}2\pi_{i}(1-\pi_{i})\,\alpha_{ik}\,\eta_{i}^{\mathrm{sig}}(k)\,\Gamma_{i}(k).$
Pick any $t_{\mathrm{src},i}\in(0,\mu_{i}^{\mathrm{truth}})$. Then the main-text asymptotic informed-truthfulness theorem holds exactly as stated: truthful (effort $e_{i}{=}1$, claim-wise truthful reporting) weakly dominates every implementable deviation and strictly dominates any uninformed deviation; inclusion under truthful converges to one. Proof. For each $k$, $\eta_{i}(k)$ is maximized at $(q_{1k},q_{0k})=(1,0)$; A3 ensures the weighted mean is positive; concentration is unchanged. ∎
Asymptotic threshold choice.
If pre-filtering ensures $\eta_{i}^{\mathrm{sig}}(k)\geq\eta_{\min}>0$ on spoken claims and $\overline{\alpha}_{i}:=\tfrac{1}{K}\sum_{k}\alpha_{ik}$ is observable, a conservative choice
$t_{\mathrm{src},i}\ \in\ \Bigl(0,\ \overline{\alpha}_{i}\,\eta_{\min}\,\gamma_{\min}\Bigr)$
guarantees $t_{\mathrm{src},i}<\mu_{i}^{\mathrm{truth}}$ and hence asymptotic inclusion under A3.
Strong truthfulness: affine inclusion (unchanged in spirit).
With $\Pr(\mathrm{include}\ i\mid\widehat{w}_{i})=a+b\,\widehat{w}_{i}$ (with $b>0$ and $a,b$ chosen so the probability is well-defined), the main-text affine strong-truthfulness theorem holds after replacing $\mu_{i}^{\mathrm{truth}}$ by the heterogeneous $\mu_{i}^{\mathrm{truth}}$ above. In particular, truthful strictly dominates if $v_{i}b\,\mu_{i}^{\mathrm{truth}}>c_{i}$. Proof. Linearity in $\mathbb{E}[\widehat{w}_{i}]$ and $\eta_{i}(k)\leq\eta_{i}^{\mathrm{sig}}(k)$ with equality only for truthful reporting establish a strict gap when $e_{i}{=}1$; for $e_{i}{=}0$ the mean score is 0. ∎
Strong truthfulness: hard threshold (what changes).
To uniformly penalize reporting deviations that flip at least a $\varphi_{\min}\in(0,\tfrac{1}{2}]$ fraction of spoken claims, we use a deterministic per-claim weight floor. Assume
$\alpha_{ik}\ \geq\ \alpha_{\min}\ >0,\qquad\eta_{i}^{\mathrm{sig}}(k)\ \geq\ \eta_{\min}\ >0\quad\text{for all }k,$
and define the per-side buffer
$\underline{g}_{i}\ :=\ \varphi_{\min}\,\alpha_{\min}\,\eta_{\min}\,\gamma_{\min}\ >0\quad\text{(half the truthful--deviation mean separation)}.$
Placing $t_{\mathrm{src},i}$ at the midpoint between the truthful mean and the worst such deviation yields this per-side buffer $\underline{g}_{i}$. With the same $3/K$ bounded-differences constant, the misclassification probability is at most $\exp(-2K\underline{g}_{i}^{2}/9)$, so for sufficiently large $K$ truthful yields strictly higher expected utility than any significant deviation. Uninformed deviations have mean 0 and are strictly dominated when $v_{i}>c_{i}$. Proof. Let $S$ be the set of flipped claims, with $|S|/K\geq\varphi_{\min}$. For $k\in S$, flipping maps $\eta_{i}^{\mathrm{truth}}(k)=\eta_{i}^{\mathrm{sig}}(k)\geq 0$ to $\eta_{i}^{\mathrm{dev}}(k)\leq 0$, so $\eta_{i}^{\mathrm{truth}}(k)-\eta_{i}^{\mathrm{dev}}(k)\geq\eta_{i}^{\mathrm{sig}}(k)\geq\eta_{\min}$. Using A3 and $\alpha_{ik}\geq\alpha_{\min}$ on $S$ gives a total expected-score separation of at least $2\underline{g}_{i}$, hence per-side buffer $\underline{g}_{i}$. Concentration then yields the utility separation as in Theorem 3.3. ∎
Finite-$K$ $\varepsilon$-informed truthfulness (what changes).
With the midpoint threshold (or any placement leaving a per-side buffer $\geq\underline{g}_{i}$), define $m_{i}:=\min\{\underline{g}_{i},\ t_{\mathrm{src},i}\}$. The main-text finite-$K$ guarantee holds with $\underline{g}_{i}$ and $m_{i}$ so defined:
[…]
Proof. Combine the per-side buffer $\underline{g}_{i}$ with the bounded-differences constant $3/K$ and apply Theorem E as in Theorem 3.4, replacing the homogeneous margin by $\underline{g}_{i}$ and $m_{i}$. ∎
Appendix I Experimental details
I.1 Data processing
@@
Natural Questions (NQ). Starting from the dev set, we filter for questions whose long-form answer has at least 100 words and 4 sentences. For clean supervision when constructing truthful paraphrases, we apply two LLM checks per item: (1) the short answer directly and correctly answers the question (not evasive or off-topic), and (2) that short answer is fully supported by the long answer. We retain only items that pass both checks and uniformly sample 300 queries. The long answer serves as the held-out gold reference answer.
ClashEval. We stratify by the six domains and sample an equal number of queries per domain (300 total). The dataset’s provided context serves as the held-out gold reference answer.
@@
For NQ, we first elicit from an LLM a plausible but incorrect short answer. We then expand this wrong answer into two non-truthful documents using fixed templates: a deceptive page (expository write-up consistent with the wrong answer) and an adversarial page (same narrative plus instruction-hijacking patterns). Prompts appear in App. I.5. For ClashEval, we use the benchmark’s provided perturbed answer (answer_mod) as the wrong narrative and apply the same two templates.
I.2 Metrics
@@
In all our experiments, LLM judges are run using gemini-2.5-flash (Comanici et al., 2025) to make the results comparable. We provide the detail prompts in I.7.
I.3 Results on Average Scores and Coherency
We first present the coherency results for the main experimental setting that is omitted in the mian text. Our method (TTS) produces summaries that are consistently more textually similar to the ground truth reference answers.
@@ I.3 Results on Average Scores and Coherency
Initial Synthesis 0.371 0.230 7.96 0.305 0.156 5.37
Majority Prompt 0.378 0.236 8.34 0.331 0.171 6.57
Majority Claims 0.367 0.216 7.36 0.303 0.152 5.20
Our Method (TTS) 0.478 0.327 14.41 0.350 0.202 8.66
@@
Table 4: Average source reliability scores ($w_{i}$) for the main experimental setting.
@@
truthful_1 0.1021 0.0876
truthful_2 0.0985 0.0828
truthful_3 0.1010 0.0890
partial 0.0402 0.0504
adversarial 0.0204 0.0258
deceptive 0.0006 0.0045
I.4 Case study: resisting coordinated, uninformative behavior.
To highlight the robustness of our method against coordinated, uninformative strategies—a canonical failure mode for consensus-based rules—we conducted a test in the ClashEval dataset involving two truthful sources, one adversarial source, and four “uninformative" sources programmed to disagree with every claim. This creates a coordinated, low-effort bloc designed to distort any mechanism based on simple agreement.
To highlight the advantage of our multi-task peer prediction scoring rule, we compare against a baseline majority scoring rule, which, to make the comparison fair, is also constructed also using leave-one-out and claim-level stances. Essentially the only difference from our mechanism is that instead of using our scoring rule (Sec. 3.1), it uses a simple majority scoring rule: $\sigma_{i}=1/K\sum_{k}\mathds{1}(r_{ik}=\text{mode}(r_{jk},\forall j))$. As shown in Result 1, traditional “majority-based" rules based on prose-level or filtering majority claims significantly underperform our approach, so we don’t include them for analysis here.
@@
truthful_1 (Truthful) 0.0226 -0.2776
truthful_2 (Truthful) 0.0209 -0.1720
uninformative_1 0.0003 0.9584
uninformative_2 0.0008 0.9660
uninformative_3 0.0006 0.9475
uninformative_4 0.0001 0.9760
adversarial -0.0001 0.1356
@@
Baseline (All Sources) 50.7% 18.4% 27.0% 6/299 (2.0%) 3
Majority-based Scoring Rule 35.8% 6.9% 11.6% 56/299 (18.7%) 86
TTS (LOO Filter) 89.2% 25.4% 39.5% 225/299 (75.3%) 43
I.5 Prompts used for generating source pool
@@
Partial (but truthful) Source
@@
For the experiment in Appendix I.4, we used four uninformative sources designed to consistently contradict any claim. The following template was used for all four, with minor variations.
@@
I.6 Prompts Used for the TTS Pipeline
@@
I.7 Prompts Used for Evaluation
@@
I.8 Ablations on model usage
@@
We evaluate TTS on 300-sample subsets from two standard information-seeking benchmarks that provide both a concise short answer and a comprehensive long-form answer for each query: Natural Questions (NQ) (Kwiatkowski et al., 2019), which pairs Google queries with annotated Wikipedia answers, and ClashEval (Wu et al., 2024), which covers six topical domains (news, names, locations, years, drugs, records). For each query, we use the long-form answer as ground truth to construct a six-document source pool from the reference answer. This pool contains four reliable sources (three high-fidelity paraphrases and one concise summary) and two unreliable sources that presents a wrong answer (one deceptive, presenting plausible but false information; one adversarial, containing prompt-injection text).

### v2

@@ Appendix G Practical Notes and Scaling for Finite-K Guarantees
Appendix H Experimental details
H.1 Data processing
@@
Natural Questions (NQ). Starting from the dev set, we filter for questions whose long-form answer has at least 100 words and 4 sentences. For clean supervision when constructing truthful paraphrases, we apply two LLM checks per item: (1) the short answer directly and correctly answers the question (not evasive or off-topic), and (2) that short answer is fully supported by the long answer. After the prefilter, we keep all samples, which gives a total of 1131 queries. The long answer serves as the held-out gold reference answer.
ClashEval. We use all unique queries in the dataset. For each query, we randomly pick a mod_degree for generating the deceptive and adversarial example. There’s a total of 1299 queries. The dataset’s provided context serves as the held-out gold reference answer.
@@
For NQ, we first elicit from an LLM a plausible but incorrect short answer. We then expand this wrong answer into two non-truthful documents using fixed templates: a deceptive page (expository write-up consistent with the wrong answer) and an adversarial page (same narrative plus instruction-hijacking patterns). Prompts appear in App. H.6. For ClashEval, we use the benchmark’s provided perturbed answer (answer_mod) as the wrong narrative and apply the same two templates.
H.2 Metrics
@@
In all our experiments, LLM judges are run using gemini-2.5-flash [Comanici et al., 2025] to make the results comparable. We provide the detail prompts in H.8. The code is available at https://github.com/jeffjiang1204/incentive-aligned-multi-source-LLM-summaries.git.
H.3 Results on Average Scores and Coherency
We first present the coherency results for the main experimental setting that is omitted in the mian text. Our method (TTS) produces summaries that are consistently more textually similar to the ground truth reference answers. As mentioned in the main text, all experimental settings here are conducted under the computation-efficiency-optimized A/B group variant described in Sec. 3.5.
@@ H.3 Results on Average Scores and Coherency
Initial Synthesis 0.375 0.238 8.20 0.330 0.166 5.75
Majority Prompt 0.385 0.247 8.56 0.355 0.183 6.77
Majority Claims 0.381 0.228 7.98 0.324 0.164 5.35
Our Method (TTS) 0.485 0.331 13.84 0.395 0.220 9.14
@@
Table 4: Average source reliability scores ($\widehat{w}_{i}$) for the main experimental setting.
@@
truthful_1 0.1072 0.0961
truthful_2 0.1079 0.0938
truthful_3 0.1068 0.0957
truthful_4 0.1033 0.0951
low_quality 0.0394 0.0526
adversarial 0.0325 0.0207
deceptive 0.0160 0.0162
H.4 Case study: resisting coordinated, uninformative behavior.
To highlight the robustness of our method against coordinated, uninformative strategies, a canonical failure mode for consensus-based rules, we conducted a test in the ClashEval dataset involving two truthful sources, one adversarial source, and four “uninformative" sources programmed to disagree with every claim. This creates a coordinated, low-effort bloc designed to distort any mechanism based on simple agreement.
To demonstrate the specific advantage of our multi-task peer prediction scoring rule, we compare it against a baseline majority scoring rule, which, to make the comparison fair, is also constructed also using leave-one-out and claim-level stances. Essentially the only difference from our mechanism is that instead of using our scoring rule (Sec. 3.1), it uses a simple majority scoring rule: $\sigma_{i}=1/K\sum_{k}\mathds{1}(r_{ik}=\text{mode}(r_{jk},\forall j))$. As shown in Result 1, traditional “majority-based" rules based on prose-level or filtering majority claims significantly underperform our approach, so we don’t include them for analysis here.
@@
truthful_1 (Truthful) 0.0308 -0.2777
truthful_2 (Truthful) 0.0302 -0.2665
truthful_3 (Truthful) 0.0315 -0.2066
uninformative_1 0.0005 0.9292
uninformative_2 0.0004 0.9310
uninformative_3 0.0006 0.9433
uninformative_4 0.0003 0.9404
adversarial 0.0022 0.6289
@@
Baseline (All Sources) 56.9% 22.6% 32.4% 7.7% 36
Majority-based Scoring Rule 34.6% 9.6% 15.1% 9.5% 189
TTS (LOO Filter) 87.8% 28.8% 43.4% 74.2% 118
H.5 Real-world case study on conflicting web evidence
Our main experiments use synthetic-but-model-generated strategic sources. This choice is driven by the specific incentive environment we study. The paper targets a forward-looking regime with increasing adoption of LLM search systems (e.g., AI overviews), where content creators optimize for how an LLM summarizes multiple documents, not just for traditional link ranking. That ecosystem is only just emerging: traditional search still dominates, and we do not yet see large-scale, systematically optimized “LLM-facing” documents at the level we anticipate if AI overviews become a default entry point.
In such an environment, there is currently no existing benchmark that closely matches our setting: multiple sources per query, some of which are deliberately strategic (deceptive, adversarial, or uninformative) with the explicit goal of shaping an LLM-generated overview. To approximate this future regime as realistically as possible, our main experiments construct strategic sources using real user queries and human-written long-form gold answers from NaturalQuestions and ClashEval and generating truthful paraphrases, low-quality sources, and deceptive / prompt-injection sources with a strong deep-reasoning model (gemini-2.5-flash), which can write content at or above typical human quality.
This yields high-quality texts that explicitly implement the behaviors we care about, deceptive alternatives, prompt injections, and coordinated uninformative blocs, while remaining grounded in real queries and reference answers. We view this as a reasonable (and in some sense conservative) proxy for how sophisticated content authors might behave when optimizing for AI-overview visibility.
To complement this synthetic setup, we systematically searched for datasets that could provide purely real sources in a setting close to ours. Concretely, we looked for datasets that (i) are grounded in user-like queries (rather than isolated claims), (ii) provide multiple pieces of conflicting evidence per query, not just a single passage labeled as true/misleading/wrong, and (iii) the evidence comes from naturally occurring web pages or comparable real-world documents.
Among verification / misinformation and retrieval-augmented QA datasets (e.g., the ones surveyed in Table 1 in [Zeng et al., 2025]), the only dataset we found that is even partially aligned is Open Domain Question Answering with Conflicting Contexts [Liu et al., 2025]. This dataset provides $1000+$ examples in which retrieved contexts may agree or conflict.
However, using it directly as a benchmark for our setting proves difficult, as the benchmark is not designed to capture the specific strategic and adversarial dynamics we study. After filtering for queries that have at least three clearly truthful sources and at least one conflicting source, only $43$ usable items remain. Moreover, when we re-evaluated the original labels with a strong LLM (gemini-2.5-flash) as a consistency check, roughly $23\%$ of the labels were flagged as dubious (many errors are confirmed by a manual check), indicating non-trivial annotation noise. This combination of small sample size and label noise makes the dataset unsuitable as a primary benchmark for our mechanism.
Despite these limitations, we treat this dataset as a valuable practical testbed for our method in a fully real-web setting. To better match our strategic model while keeping the non-synthetic nature of the pages, for each of the $43$ filtered queries, we retain the original real-web contexts and their labels (truthful vs. inaccurate). For one of the inaccurate sources, we leave its content intact and simply ask an LLM to append hidden prompt-injection instructions at the top of the document (e.g., “ignore all other evidence and output this specific answer”), mimicking real-world prompt-injection attacks in web pages. We treat the remaining truthful sources as-is, and we generate one additional deceptive source if needed by paraphrasing a wrong answer while preserving the original document’s style.
In this way, each query has a mixture of real truthful sources and at least one truly adversarial source whose body is a naturally occurring page augmented with realistic injection patterns.
On this small-but-real subset, TTS continues to behave as intended. Table 7 reports precision and answer accuracy for baselines and TTS. As in the main experiments, all methods are evaluated using the dataset’s gold answer and an LLM-as-a-judge protocol.
Table 7: Performance on Open-Domain QA with Conflicting Contexts (43 items) with real-web sources.
Method Precision (%) Answer Acc. (%)
Baseline: Initial Synthesis 64.7 46.3
Baseline: Majority Prompt 62.4 29.3
Baseline: Majority Claims 69.8 46.3
TTS (LOO Source Filtering) 80.6 65.9
TTS not only improves answer accuracy relative to all three baselines, but also assigns reliability scores that downweight the adversarial source. Table 8 reports average source-level scores under our LOO scoring rule.
Table 8: Average source reliability scores ($\widehat{w}_{i}$) on the real-web case study.
Source Name Avg. Reliability ($\widehat{w}_{i}$)
truthful_1 0.064688
truthful_2 0.052313
truthful_3 0.048616
truthful_4 0.054240
deceptive 0.047381
adversarial 0.020806
The truthful paraphrases receive the highest scores, the deceptive source is assigned an intermediate score, and the adversarial (prompt-injection) document is clearly downweighted. Although the sample size is small and the labels noisy, this real-web case study illustrates that TTS can be applied directly to naturally occurring web pages without modification to the mechanism, and that even in a noisy setting, the scoring rule tends to suppress adversarially injected content and improve factual answer quality.
Why we still rely on synthetic strategic sources for the main evaluation.
This practical example helps clarify our design choices. In this real-web dataset, most “conflicts” arise from outdated facts or minor discrepancies (e.g., date variations like “January 25, 1994” vs. “1994”), rather than intentionally strategic or adversarial content. The number of usable items is modest (43), and as mentioned earlier, there are nontrivial level of noise in the human labels, which makes it unsuitable as a primary benchmark.
By contrast, our synthetic setup lets us precisely instantiate the strategic, adversarial, and uninformative behaviors that are central to our mechanism design, and still grounds everything in real user queries and human gold answers from established benchmarks.
Taken together, we view the paper’s empirical contributions as:
  1. 1.
    controlled, forward-looking experiments on large synthetic-but-model-generated source pools (full filtered splits of ClashEval and NaturalQuestions, 1299 and 1131 queries respectively), and
  2. 2.
    a small real-web case study demonstrating that the same mechanism continues to work when applied to naturally occurring conflicting evidence.
Our goal is thus to provide a mechanism and methodology for incentive-aligned multi-source summarization, supported by both controlled simulations and a realistic stress test, rather than to propose a definitive benchmark on existing web data.
H.6 Prompts used for generating source pool
@@
Low Quality (but truthful) Source
To do this, you must omit more than 40%
The output should be a brief, standalone answer.
QUESTION:
query
REFERENCE PASSAGE:
answer
CONCISE SUMMARY:
Faithful Source 4: Blog Post
You are a blogger writing a blog post.
Your task is to rewrite the 'REFERENCE PASSAGE' as a self-contained snippet from a larger blog post.
REFERENCE PASSAGE:
answer
REWRITTEN ARTICLE SNIPPET:
Low Quality (but truthful) Source
@@
For the experiment in Appendix H.4, we used four uninformative sources designed to consistently contradict any claim. The following template was used for all four, with minor variations.
@@
H.7 Prompts Used for the TTS Pipeline
@@
H.8 Prompts Used for Evaluation
@@
H.9 Ablations on model usage
@@
We evaluate TTS on 300-sample subsets from two standard information-seeking benchmarks that provide both a concise short answer and a comprehensive long-form answer for each query: Natural Questions (NQ) [Kwiatkowski et al., 2019], which pairs Google queries with annotated Wikipedia answers, and ClashEval [Wu et al., 2024], which covers six topical domains (news, names, locations, years, drugs, records). For each query, we use the long-form answer as ground truth to construct a six-document source pool from the reference answer. This pool contains four reliable sources (three high-fidelity paraphrases and one concise summary) and two unreliable sources that presents a wrong answer (one deceptive, presenting plausible but false information; one adversarial, containing prompt-injection text).